(COMPUTE AND INTERPRET CONFIDENCE INTERVAL ESTIMATES)

An Internet service provider sampled 540 customers and found that 75 of them experienced an interruption in high-speed service during the previous month. Construct a 90% confidence interval for the proportion of all customers who experienced an interruption. Find the critical values, find E , the margin of error, then compute and interpret your interval estimate with a full sentence.

Critical value: (draw, label & shade)

Margin of error, E :

Confidence Interval:
Answer #1

Date: a I solution! Given that, n- 540 x=75 Pet75 -0.139 540 point estimate = 6.139 |- -1-0-139 = 0.86|| At 9o olo confidence